It takes 4 minutes to fill a tank with water. The tank has a hole, however, through which water drains. In 16 minutes the tank can go from full to empty. How long will it take to fill the tank if the hole is not repaired?
Answer by stanbon(57347)

It takes 4 minutes to fill a tank with water. The tank has a hole, however, through which water drains. In 16 minutes the tank can go from full to empty. How long will it take to fill the tank if the hole is not repaired?
-------------------------------
Fill rate = 1/4 job/min
---
Leak rate = 1/16 job/min
-------------------------------
Together rate = 1/x job/min
==============================
Equation:
rate - rate = together rate
1/4 - 1/16 = 1/x
---
Multiply thru by 16x to get:
4x - x = 16
3x = 16
x = 5 1/3 minutes (time to fill the tank when fill and leak are active.
================================
Cheers,
Stan H.
================